OpenStudy (jessicawade):

check my answers, polynomial identities and Pythagorean triples. will medal

OpenStudy (jessicawade):

I'm working on a project for school, but I really don't get this at all :/ The problem is : - Use the polynomial identity below you help you create a list of 10 pythagorean triples: (x^2+y^2)^2 = (x^2-y^2)^2 + (2xy)^2 Hint #1: c^2= a^2+b^2 Hint #2: pick 2 positive integers x a y, where x > y Can someone explain to me how to do this? I'm not asking to do it for me, just to explain & maybe help me with one. Thanks!

OpenStudy (jessicawade):

1.x=4 y=1 (8,15,17) 2.x=3 y=1 (5,6,7) 3.x=2 y=1 (3,4,5) 4. x=5 y=1 (10,25,27) 5.=6 y=1 (12,35,37)

OpenStudy (jessicawade):

@StudyGurl14